Human1616:

The points (3, 24) and (7, 56) represent points of a function where y, the number of photographs, varies directly with x, the number of pages in an album. Which statement describes another point on the graph of this function?

snowflake0531:

Firstly, you would have to find the equation of this. Then, you can just state a random point, but x has to be positive. Do you know how to find the slope of this?

Human1616:

is that where i put y2-y1 over x2-x1

snowflake0531:

Yes, that's correct! Do you know how to do that?

Human1616:

so i would start with 56-24 witch equals 32 then i would subtract 3 from 7 and get 4

snowflake0531:

Yes!, So then it would be 32/4, which is 8 So, there's our slope! But then, we have to find the y-intercept to know the whole equation Choose one of the two coordinates, either (3,24) or (7,56) and plug the x and the y, and also the slope, into the equation y=mx+b, and solve for b

Human1616:

y=3x+8 is that correct or is it y=24x+8

snowflake0531:

How did you do this?

snowflake0531:

Nevermind, So, I would use the first coordinate point, (3,24) and the slope is 8 So, 24 = 3(8) +b 24=24+b so, b=0 Therefore, the equation of the line would be y= 8x Ye?

Human1616:

ye that makes sense

snowflake0531:

Okay, so "Which statement describes another point on the graph of this function?" and also "where y, the number of photographs, varies directly with x, the number of pages in an album" Choose a random x value

Human1616:

8 so it is an easy number to multiply by

snowflake0531:

Okay, and plugging in 8 as the value of x in y=8x, what is y? I suggest multiplying 8 by 8

Human1616:

64

snowflake0531:

Okay, so y=64 So, when x is 8, y is 64 What does this mean? *referencing back to "where y, the number of photographs, varies directly with x, the number of pages in an album"

Human1616:

64 photographs to 8 pages?

snowflake0531:

Yes, but I just reread the question It seems like this is a multiple choice question... oops. What are the choices.

Human1616:

i figured it out by looking at the answer choices and got the right one thank you for your help

snowflake0531:

Yw!~